OpenStudy (anonymous):

write y = -.566021616 (x - 6) ^2 + 3.7 in general form?

OpenStudy (anonymous):

(x-6)^2=x^2-12x+36 as (a-b)^2= a^2-2ab+b^2 multiply it wd -.566021616 den simplfy...